OpenStudy (anonymous):

Prove: Let V(x) and D(x) be polynomials, then there exists exactly one polynomial Q(x) and exactly one polynomial R(x) where: V(x) = Q(x)*D(x)+R(x) and (R(x)=0 or degree[R(x)]

OpenStudy (anonymous):

rewrite like V(x)/Q(x) = D(x) + R(x)/Q(x)

OpenStudy (anonymous):

yes, but then? :D

OpenStudy (anonymous):

*

OpenStudy (anonymous):

hmm, not sure. I though we were going to be able to use the remainder theorem in a pretty simple way to show this, but it's not coming to me... I'll have to think about it some more, maybe someone can jump in if they have an idea..

OpenStudy (anonymous):

Well, it is quite possible related to the remainder theorem. It is stated in my textbook, just above the remainder theorem. But no proof is given so I find it hard to understand.

OpenStudy (amistre64):

numbers are called polynomials of degree 0, so it seems to be a more general approach to the thrm

OpenStudy (anonymous):

Unless the number is 0. Then it is a polynomial of degree "undefined" or degree "-infinity"

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

think of polynomials as encrypted "numbers".

OpenStudy (amistre64):

since they are functions, then for any value of x, there is a distinct value for y that results

OpenStudy (anonymous):

I guess 12/3 = 4 and R=0. degree of R(x) has to be less than degree of D(x) or R(x) has to be zero. So I guess that's a good start, since you can't give any other Q(x) than 4 where R=0

OpenStudy (anonymous):

But it's not a proof...

OpenStudy (amistre64):

the proof would start out like this most likely let\[P(x) = p_nx^n+p_{n-1}x^{n-1}+p_{n-2}x^{n-2}+...+p_2x^2+p_1x+p_0\]let\[Q(x) = q_nx^n+q_{n-1}x^{n-1}+q_{n-2}x^{n-2}+...+q_2x^2+q_1x+q_0\]

OpenStudy (amistre64):

There is one polynomial, D(x) such that\[D(x)=\frac{p_nx^n}{q_nx^n}+...\] maybe swap out ns with as and bs to indicate the possibility of different degrees

OpenStudy (amistre64):

\[q_2x^2+q_1x+q_0~|~p_3x^3+p_2x^2+p_1x+p_0\] D(x) =\(\large \frac{p_3x^3}{q_2x^2}+\frac{}{}+\frac{}{}+\) --------------------------- q_2x^2+q_1x+q_0 | \(p_3x^3+p_2x^2+p_1x+p_0\) \(-p_3x^3\) latex needs a long hand division mode ....

OpenStudy (anonymous):

also search about Division Algorithm of polynomials

OpenStudy (anonymous):

But can you divide a polynomial of lower degree by one of higher degree? Wouldn't that disprove the theorem?

hartnn (hartnn):

guys, why are u complicating things ., showing the uniqueness is real simple. use contradiction, and prove that its not possible to have 2 r(x)'s.

OpenStudy (amistre64):

you cant divide a lower degree by a higher and still get a polynomial back

OpenStudy (amistre64):

or something like that

OpenStudy (anonymous):

Yes, doesn't that disprove the theorem?